OpenStudy (micheal1234):

Addition, Subtraction, Multiplication & Division Properties of Equality Equivalent Equations Isolate Inverse Operations 7. Solve for the variable: 𝑘𝑘 4=−17/ 2 8. Solve for the variable: 3/8𝑝=9 9. What is the solution of -3 = x + 5 ? A.-15 B. -8 C. 2 D.8

OpenStudy (micheal1234):

Anaise, and Rainbow_Factory please help

OpenStudy (micheal1234):

any idieas

OpenStudy (anonymous):

Are the these --> 𝑝 <-- supposed to be a a blank box or no because my computer is kind old..

OpenStudy (micheal1234):

sorry about that do not worry about them

OpenStudy (anonymous):

Ok

OpenStudy (micheal1234):

wait

OpenStudy (anonymous):

?

OpenStudy (micheal1234):

7.𝑝𝑝 is kk 8.𝑝is p

OpenStudy (anonymous):

I don't know about these boxes I'm having a hard time with them half of the question is blocked with the boxes

OpenStudy (micheal1234):

the first two boxes are k the lass one one is p i mess up putting them on there.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

lol

OpenStudy (micheal1234):

i know

OpenStudy (anonymous):

do u think u can copy your problem page and share it with me so there are no boxes please?

OpenStudy (micheal1234):

ok

OpenStudy (anonymous):

Thanks

OpenStudy (micheal1234):

7. Solve for the variable: kk4=−17/ 2 8. Solve for the variable: 3/8p=9 9. What is the solution of -3 = x + 5 ? A.-15 B. -8 C. 2 D.8

OpenStudy (micheal1234):